Student affairs departments at colleges provide programs and events that offer students opportunities to learn outside of tranb

Partic Retained

0

1

1

1

1

0

1

1

0

1

1

1

0

1

1

1

1

0

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

0

1

0

1

0

1

1

0

1

0

0

1

1

1

1

1

0

1

1

0

0

1

0

1

1

1

1

1

1

1

1

1

0

1

1

1

0

1

1

1

0

0

1

1

1

1

0

1

1

1

0

1

1

1

1

0

1

1

1

1

0

0

1

1

1

1

1

1

0

1

0

0

1

1

1

1

1

1

1

1

1

1

1

0

1

1

1

1

0

1

1

1

0

0

1

1

1

1

0

1

0

1

0

1

1

0

0

1

1

1

0

1

0

1

1

1

1

1

1

0

1

0

1

1

0

1

0

1

1

1

1

1

0

0

0

1

1

1

1

1

1

1

1

1

1

1

1

0

1

1

1

1

1

1

1

1

1

1

0

1

0

1

0

1

1

1

1

1

1

1

1

1

1

1

1

1

0

1

1

0

0

1

1

0

1

1

0

1

1

1

1

1

0

1

0

1

1

1

1

1

1

1

1

1

1

1

0

1

1

1

1

0

1

1

1

1

1

1

1

1

1

1

0

1

0

1

0

1

0

1

1

0

1

1

0

0

1

1

1

0

1

1

1

1

1

1

1

1

1

1

1

1

0

1

1

1

0

0

1

1

1

1

1

1

1

1

1

0

1

0

0

0

1

1

1

1

1

1

1

1

0

1

1

1

1

1

1

1

1

1

1

1

0

1

1

0

0

1

1

1

1

1

1

1

1

0

1

1

1

1

1

1

1

1

0

1

0

1

1

1

1

0

1

1

1

0

0

0

0

0

1

1

1

1

0

1

1

1

0

1

1

0

0

1

0

1

0

1

1

0

1

1

0

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

0

1

0

0

0

1

1

0

1

1

1

1

1

1

0

1

1

1

1

0

1

1

1

1

1

0

1

1

1

0

0

0

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

0

0

1

1

0

1

1

1

0

0

1

1

1

1

1

0

0

1

1

1

1

0

1

0

1

0

0

0

1

1

1

1

1

1

1

1

1

0

1

1

0

1

0

1

1

0

0

0

1

1

0

1

0

1

1

0

1

1

1

1

0

1

1

1

1

1

1

0

1

1

1

0

0

1

1

0

1

1

1

1

1

1

1

1

1

1

0

1

1

1

1

1

0

1

1

1

1

1

1

1

1

1

0

1

1

1

1

1

1

1

1

0

1

1

1

1

1

0

1

1

1

0

1

1

0

1

0

1

0

0

1

0

1

1

1

0

0

1

1

0

0

1

1

0

1

1

1

1

1

0

1

0

0

1

1

1

1

1

1

1

1

0

1

1

1

1

1

1

1

1

1

0

0

1

1

1

1

0

1

1

1

0

1

1

1

1

1

0

0

1

1

1

0

1

0

0

0

0

1

1

0

1

0

0

1

1

1

1

0

1

1

1

1

1

1

1

0

1

0

0

1

1

1

0

1

1

1

1

1

0

0

1

1

1

1

1

1

1

1

1

0

1

1

0

0

1

1

0

1

1

1

1

1

1

1

1

1

1

0

1

1

0

1

1

1

1

1

1

1

1

1

1

1

1

1

1

1

0

1

1

1

1

1

1
Answer #1

Test natur solution Here Ho P=0.73 vs HIP >0.73 D from the given data o = 56 = 0.75973 c z= pl-p= 0.75973-0.73 = 1.828 pa (0.